Observatory Log: February 19, 2013

Doyle and I went out to the observatory tonight. It was one of the coldest nights of this season. It was also my last trip to the Paul P Feder Observatory (being a student researcher). My goal was taking pictures of the instruments including the spectrograph assembly, all of which will be used  in the procedure manual I have been working on for astronomy students at MSU-Moorhead. We found some snow on the floor of the dome. See the attached picture. It was a productive night. I will take this opportunity to thank Doyle Heden who volunteered and helped me to finish the task on time.

Snow on the floor

Posted in Observing Log

Observatory Log: January 25, 2013

Nathan and I went out to the observatory tonight. The sky was clear when we opened the dome at 7:30 pm. The ambient temperature was around 4.1 F. We initialized the telescope and mounted the spectrograph assembly at 8:20 pm. The rest of the night was uneventful, since we could not align the star on the slit. We changed the focus, but it did not help much. I changed the alignment of the spectrograph. First the the camera was facing toward the cable outlet. Later, I oriented the spectrograph parallel to the fork. But nothing worked. We might be out of focus. Temperature dropped to -0.2 F when we called it a night at 9:30 pm.

Posted in Observing Log

Observing Log: January 21, 2013

Matt and I went out to the observatory tonight. Both seeing and transparency were below average. Matt spent about an hour following the instructions for the start up procedure at the observatory. We mounted the spectrograph assembly around 8:20 pm and acquired data of Ne spectra using both the low and high resolution mode of the spectrograph. I am looking forward to spend some time on those two files and try to get some information of wavelength versus pixels. The ambient temperature was 7.4 F and the dome temperature was 6.2 F, when we called it a night.

Posted in Observing Log

Observing Log: Decemeber 31, 2012

Matt and I went out to the observatory tonight. Transparency was below average. We mounted the spectrograph around 8 pm (after initializing the telescope to the set star). As we discussed earlier, we changed the focus of the telescope slowly, but both the motor and the computer stopped responding when the focus was around 3200 [Task manager was on the desktop all the time]. The ambient temperature was 3.1 F.  The star was on the slit, when I checked the focus at 3100.

Posted in Observing Log

Observing Log: September 15, 2012

Aaron and I went out to the observatory tonight. The sky was clear and transparency was average. We updated the license information for newly purchased MaxIM DL 5 software. We encountered an error message on the intelligent filter wheel when we were using the MaxIM DL 5 software.The error message was :”Error=2″, which implies that the software did not identify that there were four filters already installed in the filter wheel chamber.We switched to the older version of MaxIM DL software afterwards. We had an issue with the dome too. The telescope slewed to 270.4 and the DFM-TCS software read that to be its zenith position.The computer froze when we tried to fix this using the TCS software. We slewed the telescope to its zenith position manually.We had to reboot the computer. We tried to initialize one more time, but it did not work as the dome did not move to the proper position. Before we shut down the system, the telescope was pointing in the direction that it was supposed to, but the field view was covered by the dome.We decided to shut down the system following the typed up procedures and called it a day.

Posted in Observing Log

Observing Log, Aug 20, 2012

Dr. Craig, Dr. Cabanela, Aaron and I went out to the observatory tonight. Aaron and I spent the first half of the night adjusting the set star, which was supposed to appear on the illuminated slit.We got to see the star, but it was out of focus. When we started changing the focus, we completely lost track of it.

Tagged with:
Posted in Observing Log

Observing Log, Aug 13, 2012

Dr. Craig and I went to the observatory tonight. The sky was partially cloudy. Transperancy was moderately good. We mounted the spectrograph and spent some time setting the focus. The focus was 3256 at an ambient temperature of 68 F. Again, always cross check the sky clock with FM weather website before you depart.

Tagged with:
Posted in Observing Log